OpenStudy (anonymous):

Ok guys help out again find the next number in a sequence 1,8,16,25. I'm confused on how u do this

OpenStudy (anonymous):

1+7=8 8+8=16 16+9=25 Every number is found by adding one more than what was added to the previous number. 25+10=35 35+11=46 so on...

OpenStudy (anonymous):

Thank u. How about this one 1,-1/2,1/4

OpenStudy (anonymous):

The pattern is dividing each number by 2 1/2=1/2 .5/2=1/4 .25/2=1/8